### v1

# AlignGPT: Multi-modal Large Language Models with Adaptive Alignment Capability

Abstract
Multimodal Large Language Models (MLLMs) are widely regarded as crucial in the exploration of Artificial General Intelligence (AGI). The core of MLLMs lies in their capability to achieve cross-modal alignment. To attain this goal, current MLLMs typically follow a two-phase training paradigm: the pre-training phase and the instruction-tuning phase. Despite their success, there are shortcomings in the modeling of alignment capabilities within these models. Firstly, during the pre-training phase, the model usually assumes that all image-text pairs are uniformly aligned, but in fact the degree of alignment between different image-text pairs is inconsistent. Secondly, the instructions currently used for finetuning incorporate a variety of tasks, different tasks’s instructions usually require different levels of alignment capabilities, but previous MLLMs overlook these differentiated alignment needs. To tackle these issues, we propose a new multimodal large language model AlignGPT. In the pre-training stage, instead of treating all image-text pairs equally, we assign different levels of alignment capabilities to different image-text pairs. Then, in the instruction-tuning phase, we adaptively combine these different levels of alignment capabilities to meet the dynamic alignment needs of different instructions. Extensive experimental results show that our model achieves competitive performance on 12 benchmarks.
Figure 1: AlignGPT achieves competitive performances on a broad range of vision-language tasks compared with other generalist models. To facilitate observation, we only show the performance of MiniGPT-v2 and AlignGPT. More comprehensive data can be found in Table 3.3 and Table 3.3.
1 Introduction
Multimodal Large Language Models (MLLMs) are considered a crucial step towards achieving Artificial General Intelligence (AGI) DBLP:journals/corr/abs-2303-08774 ; DBLP:journals/corr/abs-2312-11805 ; DBLP:conf/icml/DriessXSLCIWTVY23 ; DBLP:journals/corr/abs-2309-16058 . The uniqueness of these models lies in their ability to integrate and understand various types of information, especially text and image data. In the pursuit of AGI, this cross-modal understanding and processing capability is essential, as it mimics how humans interact with the world and comprehend complex information through different senses, such as vision and language. The development of multimodal large language models not only advances the field of artificial intelligence but also provides machines with a way to process and understand information that is closer to human cognition.
Currently, MLLMs typically adhere to a unified training paradigm, which is divided into two key phases: the pre-training phase and the instruction-tuning phase DBLP:conf/nips/LiuLWL23a ; DBLP:journals/corr/abs-2304-10592 ; DBLP:journals/corr/abs-2304-14178 ; DBLP:journals/corr/abs-2308-12966 ; DBLP:journals/corr/abs-2310-09478 ; DBLP:journals/corr/abs-2311-03079 ; DBLP:conf/aaai/HuXLLCT24 ; DBLP:journals/corr/abs-2309-05519 ; DBLP:journals/corr/abs-2309-15112 . The pre-training phase concentrates on aligning images with text, aiming to train the model to understand the relevance of image contents and their respective textual descriptions. This alignment imbues the model with cross-modal comprehension abilities. The instruction-tuning phase further enhances its adaptability to specific tasks. This includes enabling the model to complete particular visual-language tasks based on given instructions, such as generating textual descriptions from images, answering questions related to images, or even performing complex reasoning based on both text and images. This training paradigm equips multimodal pre-trained models with not only fundamental cross-modal understanding but also the flexibility to adapt to diverse task demands.
Although current MLLMs have made great progress, the modeling of alignment capabilities during their pre-training and instruction-tuning phases is still insufficient for the following reasons:
  • The degree of alignment is inconsistent between different image-text pairs: During the pre-training phase, the model typically operates on a key assumption that all image-text pairings are consistently aligned. However, in practical scenarios, the degree of alignment in image-text pairings is not always uniform: in some image-text pairs, the text may describe the whole image while in others the text only describes a part of the image. If these differences are not differentiated during the pre-training phase, it could lead to a misunderstanding of the image-text alignment relationships in the learning process;
  • The instructions for different tasks require different levels of alignment capabilities: The instructions currently used for finetuning incorporate a variety of tasks. Some of these tasks, like image captioning DBLP:conf/icml/XuBKCCSZB15 , rely more on global image-text alignment capabilities. In contrast, other tasks, such as visual question answering (VQA) DBLP:conf/iccv/AntolALMBZP15 , typically require the model to answer questions based on specific parts of the image. This necessitates not only global image-text alignment but also local image-text alignment capabilities. Thus, the instructions of different tasks demand different levels of alignment capabilities.
To effectively enhance the alignment capabilities during the pre-training and instruction-tuning phases, we propose a new multimodal large language model called AlignGPT. In the pre-training phase, we introduce a new paradigm with controllable alignment levels, which does not treat all image-text pairs equally; instead, it assigns different levels of alignment capability to different pairs. This is achieved through CLIP scores DBLP:conf/icml/RadfordKHRGASAM21 , where the model categorizes image-text pairs into different alignment levels based on their CLIP scores. A higher alignment level indicates that the text contains more comprehensive information about the image DBLP:journals/corr/abs-2111-02114 ; DBLP:conf/icvgip/GroverMG22 . Subsequently, we utilize these alignment levels as control signals to address the issue of varying degrees of alignment in image-text pairings. During the instruction-tuning phase, we first transform these different levels of alignment capabilities obtained by pre-training into global and local alignment capabilities based on the strength of the alignment levels. Then, we not only assign global alignment capabilities to the instructions of each task, but also adaptively configure different local alignment capabilities according to the alignment needs of each instruction. The broad range of tests conducted demonstrates that our model achieves competitive performance across 12 benchmarks, as shown in Figure 1.
Our contribution can be summarized as follows: (1) We propose a new multi-modal large language model AlignGPT to elevate and empower the alignment capabilities of MLLMs; (2) We propose a novel alignment strategy that generates different levels of alignment capabilities in the pre-training stage, and then adaptively combines these alignment capabilities in the instruction-tuning stage to meet the alignment needs of different instructions; (3) We conduct evaluations across multiple academic benchmarks and multimodal instruction-following benchmarks. Extensive experimental results show that our proposed AlignGPT achieves competitive performance.
2 Related Work
In this section, we review the existing studies on large language models and visual language models.
Large Language Models.
In the field of natural language processing, BERT DBLP:conf/naacl/DevlinCLT19 and GPT-2 radford2019language , as pioneering large pre-trained language models, marked a significant breakthrough in this technological direction. Their training on vast web text datasets demonstrated unprecedented language understanding and generation capabilities. Subsequently, the launch of GPT-3 DBLP:conf/nips/BrownMRSKDNSSAA20 further accelerated the development of this field, with its large model parameters and extensive training datasets showcasing exceptional abilities in few-shot learning, significantly enhancing task adaptability and flexibility. Following this, the introduction of InstructGPT and ChatGPT DBLP:conf/nips/Ouyang0JAWMZASR22 focused on optimizing the efficiency and naturalness of interactions between models and humans, where InstructGPT enhanced the capability to execute precise instructions, and ChatGPT improved the conversational experience, making these models more fluent in human-computer communication. Meanwhile, as large language model (LLM) technology continued to evolve, emerging models like LLaMA DBLP:journals/corr/abs-2302-13971 and GLM du2022glm began to make their mark. To equip these models with the ability to respond to human instructions similar to ChatGPT, research teams finetune LLaMA and GLM using high-quality instruction datasets, thereby further enhancing its capability to follow instructions, with representative projects such as Alpaca taori2023alpaca , Vicuna chiang2023vicuna , and ChatGLM DBLP:conf/iclr/ZengLDWL0YXZXTM23 .
Although these models have made significant progress in interacting with humans through language, we recognize that human understanding and processing of complex information relies not only on language but also critically on visual and other sensory inputs. The observation has driven us to further explore more comprehensive visual-language models in order to more accurately simulate complex interactions between humans and the real world.
Visual Language Models.
In recent years, multimodal large language models (MLLMs) have garnered increasing attention. The core of MLLMs lies in their ability to achieve cross-modal understanding and generalization. Most current models, such as LLaVA DBLP:conf/nips/LiuLWL23a , MiniGPT-4 DBLP:journals/corr/abs-2304-10592 , mPLUG-Owl DBLP:journals/corr/abs-2304-14178 , Qwen-VL DBLP:journals/corr/abs-2308-12966 , MiniGPT-v2 DBLP:journals/corr/abs-2310-09478 , NExT-GPT DBLP:journals/corr/abs-2309-05519 , InternLM-XComposer DBLP:journals/corr/abs-2309-15112 , CogVLM DBLP:journals/corr/abs-2311-03079 , and MM1 DBLP:journals/corr/abs-2403-09611 , utilize a standard training framework consisting of two primary phases: pre-training and instruction-tuning. In the pre-training phase, the model utilizes image caption data to establish a rich understanding of cross-modal semantic knowledge. This training phase enables the model to comprehend and capture the correlation between images and text, establishing a solid foundation for subsequent stage. In the instruction-tuning phase, the model receives specific task instructions to optimize its performance on that task. Through this instruction-tuning phase, the model can further refine its understanding to execute specific tasks, enabling it to flexibly and accurately address various task requirements in practical applications.
Although achieving good results, the current MLLMs overlook two critical factors: first, the degree of alignment between different image-text pairs is inconsistent during the pre-training phase; second, the instructions for different tasks require different levels of alignment capabilities during the instruction-tuning phase. Therefore, the modeling of alignment capabilities in these models remains inadequate. To this end, we propose a new multimodal large language model AlignGPT to effectively enhance the alignment capabilities of MLLMs.
3 Methodology
In this section, we initially present the fundamental structure of the visual-language model AlignGPT, followed by a demonstration of how to enhance the alignment capability of the model during both the pre-training and instruction-tuning stages.
Figure 2: The architecture of AlignGPT.
3.1 Architecture
AlignGPT consists of four components: a visual backbone, a linear projection layer, a large language model, and an alignment module. Figure 2 provides an overview of the AlignGPT architecture and its training process. The followings are the implementation details of these components:
Visual backbone.
We utilize the pre-trained CLIP visual encoder ViT-L/14 DBLP:conf/icml/RadfordKHRGASAM21 as our visual backbone. We train the model using an image resolution of 336$\times$336.
Linear projection layer.
We adopt a linear projection layer to map the representations of images from the vector space of the vision backbone to that of the language model.
Large language model.
We choose the open-source model Vicuna chiang2023vicuna as our language model backbone, given its strong ability to follow instructions effectively in various language tasks.
Alignment module.
We propose to add alignment vectors to the inputs of MLLMs to enrich their alignment capabilities. These alignment vectors are positioned ahead of the image embeddings and text embeddings. In the subsequent sections, we will elaborate on the role of the alignment vectors and the process to acquire them.
3.2 Alignment Level-aware Pre-training
As mentioned before, in the pre-training stage, the model usually assumes that all image-text pairs are uniformly aligned, and these pairs are used to train the model to comprehend the relations between images and their corresponding textual descriptions. However, the real-world scenarios show that the degree of alignment between these image-text pairs may vary considerably. Overlooking the difference could lead to a misunderstanding of the image-text alignment relations during the learning process.
Instead of treating all image-text pairs equally, we assign different levels of alignment capabilities to different image-text pairs. To achieve this, we leverage the similarity scores provided by CLIP. The motivation behind this is that CLIP is pre-trained on a large-scale image-text dataset, thus using CLIP similarity scores to determine alignment levels between images and text is relatively accurate. Image-text pairs with lower CLIP scores suggest that the text describes only part of the image’s information, whereas pairs with higher CLIP scores indicate that the text provides a more comprehensive description of the image DBLP:journals/corr/abs-2111-02114 ; DBLP:conf/icvgip/GroverMG22 . Subsequently, we use these alignment levels as control signals to train the model, enabling it to understand the alignment relations between different image-text pairs.
More precisely, we start by computing the CLIP similarities $s$ for all training image-text pairs. Then, we rank all image-text pairs based on their similarity scores. Finally, we use a bucketing technique to divide them into $N$ discrete alignment levels. The process can be represented as:
$l=bucket(s),\quad l\in\{1,2,...,N\},$ (1)
where $bucket(\cdot)$ denotes a bucketing function that assigns each pair into one of $N$ equally spaced intervals and $l$ is the alignment level of an image-text pair. In this way, image-text pairs with lower CLIP similarity scores are assigned to buckets indicative of lower alignment levels, whereas those with higher CLIP similarity scores are grouped into buckets representing higher alignment levels.
Once the alignment level of each image-text pair is determined, we can regard it as a special token to express the alignment relation between the image and its textual description. This special token is placed ahead of the image and text tokens. During the pre-training phase, in addition to learning the mapping function from the image to the text space in the linear projection layer, we also initialize this special token as an alignment vector and continuously update its representation.
3.3 Adaptive Alignment-based Instruction-tuning
Currently, the instructions used for finetuning cover various tasks such as image captioning, visual question answering, and visual grounding, etc. The instructions of these tasks place different requirements on the alignment capabilities. For example, image captioning tasks mainly rely on global alignment between images and text, while VQA and visual grounding tasks require not only global alignment but also local alignment capabilities between images and text.
To this end, we first demonstrate how to represent the global and local alignment capabilities between image-text pairs. As mentioned in Section 3.2, after the pre-training stage, we obtain $N$ alignment vectors $\{H_{1},H_{2},...,H_{N}\}$ corresponding to $N$ discrete alignment levels $\{1,2,...,N\}$. Among them, $H_{N}$ represents the highest level of alignment, i.e., $H_{N}$ indicates that the text provides very comprehensive description of an image. Here we regard it as a global alignment vector. The vectors below $H_{N}$ represent different degrees of alignment between the image and the text (i.e., $\{H_{1},H_{2},...,H_{N-1}\}$), which means the text only describes a part of the information of the image from weak to strong. Thus, we regard them as local alignment vectors of varying degrees.
Afterwards, we not only allocate global alignment capabilities to the instructions of each task, but also adaptively distribute varying degrees of local alignment capabilities based on the distinct alignment needs of each instruction. The reason behind this is that global alignment serves as the foundation for cross-modal understanding; only by mastering global alignment capabilities can a model truly focus on enhancing local alignment abilities. Specifically, in addition to the global alignment vectors, we assign different weights to the local alignment vectors via a gate network. These weights are obtained based on input instructions and image, as the input instructions greatly influence the visual regions the model should focus on. The implementation of the gate network is as follows:
$\alpha=softmax(W(H_{I}\otimes H_{T})+b),$ (2)
where $H_{I}$ and $H_{T}$ denote the vector representation of input instruction and image information, $W$ and $b$ are weight matrix and bias, $\alpha$ means the weights of local alignment vectors. Finally, we aggregate the global alignment vector and the local alignment vectors with varying weights to ensure a more precise fulfillment of alignment requirements for each instruction:
$H_{align}=H_{N}+\sum_{i=1}^{N-1}\alpha H_{i},$ (3)
where $H_{align}$ means the alignment vector for each instruction during the instruction-tuning stage.
Overall, we can view the alignment vectors obtained during the pre-training phase as foundational components, each with varying alignment capabilities. During the instruction-tuning phase, we dynamically combine these components to meet the alignment needs of different instructions.
Table 1: Performance comparison on multiple academic benchmarks. Res indicates the resolution of the input image. For the baselines, the results with are obtained by running the code released by the authors, and the other results are retrieved from DBLP:journals/corr/abs-2310-09478 ; DBLP:journals/corr/abs-2310-03744 . Best results are in bold.
{tabu}
l|l|c|cc|ccccc Method LLM Res. Sample Size   VQAv2 GQA VisWiz SQAI TextVQA
Pre-train Finetune
BLIP-2 Vicuna-13B 224 129M - 41.0 41.0 19.6 61.0 42.5
InstructBLIP Vicuna-7B 224 129M 1.2M - 49.2 34.5 60.5 50.1
InstructBLIP Vicuna-13B 224 129M 1.2M - 49.5 33.4 63.1 50.7
Shikra Vicuna-13B 224 600K 5.5M 77.4 - - - -
IDEFICS-9B LLaMA-7B 224 353M 1M 50.9 38.4 35.5 - 25.9
IDEFICS-80B LLaMA-65B 224 353M 1M 60.0 45.2 36.0 - 30.9
MiniGPT-v2 LLaMA2-7B 448 - - 74.6 60.3 32.9 60.9 28.0
Qwen-VL Qwen-7B 448 1.4B 50M 78.8 59.3 35.2 67.1 63.8
Qwen-VL-Chat Qwen-7B 448 1.4B 50M 78.2 57.5 38.9 68.2 61.5
LLaVA-1.5 Vicuna-7B 336 558K 665K 78.5 62.0 50.0 66.8 58.2
LLaVA-1.5 Vicuna-13B 336 558K 665K 80.0 63.3 53.6 71.6 61.3
AlignGPT Vicuna-7B 336 558K 665K 79.1 62.9 54.2 68.5 58.4
AlignGPT Vicuna-13B 336 558K 665K 80.0 63.6 56.4 70.3 60.2
Table 2: Results on multimodal instruction-following benchmarks. For the baseline methods, the results on the SEEDI dataset are obtained from DBLP:journals/corr/abs-2311-12793 , and the other results are retrieved from DBLP:journals/corr/abs-2310-03744 .
{tabu}
l|l|c|ccccccc
Method LLM Res. POPE MME MMB MMBCN SEEDI LLaVAW MM-Vet
BLIP-2 Vicuna-13B 224 85.3 1293.8 - - 46.4 38.1 22.4
InstructBLIP Vicuna-7B 224 - - 36.0 23.7 53.4 60.9 26.2
InstructBLIP Vicuna-13B 224 78.9 1212.8 - - - 58.2 25.6
Shikra Vicuna-13B 224 - - 58.8 - - - -
IDEFICS-9B LLaMA-7B 224 - - 48.2 25.2 - - -
IDEFICS-80B LLaMA-65B 224 - - 54.5 38.1 - - -
MiniGPT-v2 LLaMA2-7B 448 85.1 1332.1 43.1 29.1 52.3 - -
Qwen-VL Qwen-7B 448 - - 38.2 7.4 56.3 - -
Qwen-VL-Chat Qwen-7B 448 - 1487.5 60.6 56.7 58.2 - -
LLaVA-1.5 Vicuna-7B 336 85.9 1510.7 64.3 58.3 66.2 63.4 30.5
LLaVA-1.5 Vicuna-13B 336 85.9 1531.3 67.7 63.6 68.2 70.7 35.4
AlignGPT Vicuna-7B 336 86.0 1527.4 67.3 59.9 66.5 68.4 30.8
AlignGPT Vicuna-13B 336 86.2 1572.0 69.5 63.7 67.8 75.2 35.6
4 Experiments
4.1 Experimental Settings
Datasets.
For a fair comparison, we use the same pre-training and instruction dataset as the LLaVA-1.5 DBLP:journals/corr/abs-2310-03744 . It mainly includes 558K caption pairs for modality alignment and 665K single- or multi-round conversations for instruction-tuning. Besides, we evaluate AlignGPT on a range of academic visual question answering (VQA) tasks and recent benchmarks designed specifically for MLLMs. This evaluation spans 12 benchmarks, including VQAV2 DBLP:conf/cvpr/GoyalKSBP17 , GQA DBLP:conf/cvpr/Gurari0SGLGLB18 , VizWiz DBLP:conf/cvpr/Gurari0SGLGLB18 , SQAI (ScienceQA-IMG) DBLP:conf/nips/LuMX0CZTCK22 , TextVQA DBLP:conf/cvpr/SinghNSJCBPR19 , POPE DBLP:conf/emnlp/LiDZWZW23 , MME DBLP:journals/corr/abs-2306-13394 , MMB (MMBench), MMBCN (MMBench-Chinese) DBLP:journals/corr/abs-2307-06281 , SEEDI (SEED-Bench-IMG) DBLP:journals/corr/abs-2307-16125 , LLaVAW (LLaVA-Bench-in-the-Wild) DBLP:conf/nips/LiuLWL23a , and MM-Vet DBLP:journals/corr/abs-2308-02490 datasets.
Implementation Details.
We adopt a ViT dosovitskiy2021an model pre-trained with CLIP DBLP:conf/icml/RadfordKHRGASAM21 as a vision encoder to effectively process visual inputs. On the language side, Vicuna chiang2023vicuna is utilized to handle multimodal features, ensuring a cohesive integration of text and visual data. In the pre-training phase, both the visual backbone and the large language model of AlignGPT remain frozen, with only the parameters of the linear projection layer and alignment vectors being trained. During instruction-tuning phase, we freeze the linear projection layer, alignment vectors, and visual backbone, while adjusting the parameters of the large language model and the gate network. The global batch sizes for the two phases are set at 256 and 128 respectively, with DeepSpeed DBLP:conf/sc/RajbhandariRRH20 using ZeRO2 and ZeRO3 strategies accordingly. Regarding our training methodology, we conduct a single epoch of optimization for all models using the AdamW loshchilov2018decoupled optimizer coupled with a cosine learning schedule. Moreover, we initiate pre-training and instruction-tuning with learning rates of 1e-3 and 2e-5, respectively. The framework is trained on 8 A800 GPUs with 80GB memory.
4.2 Compared Methods
We chose a diverse set of representative multimodal large language models (MLLMs) as our baselines, including BLIP-2 DBLP:conf/icml/0008LSH23 , InstructBLIP DBLP:conf/nips/Dai0LTZW0FH23 , Shikra DBLP:journals/corr/abs-2306-15195 , IDEFICS DBLP:conf/nips/LaurenconSTBSLW23 , MiniGPT-v2 DBLP:journals/corr/abs-2310-09478 , Qwen-VL DBLP:journals/corr/abs-2308-12966 , Qwen-VL-Chat DBLP:journals/corr/abs-2308-12966 , and LLaVA-1.5 DBLP:journals/corr/abs-2310-03744 .
5 Results and Analysis
5.1 Main Results
Visual Question Answering.
We evaluate AlignGPT using five popular academic benchmarks, as detailed in Table 3.3. Despite using less training data, the AlignGPT-7B demonstrates competitive performance, surpassing other generalist models including InstructBLIP-13B, Shikra-13B, and IDEFICS-80B on most datasets, except for LLaVA-1.5-13B. These results verify the rationality of the structural design of our model. Moreover, considering that AlignGPT utilizes the same training dataset as LLaVA-1.5, it is evident that AlignGPT-7B outperforms LLaVA-1.5-7B across all evaluation datasets, and AlignGPT-13B also surpasses LLaVA-1.5-13B on the majority of datasets. This demonstrates that our approach has effectively enhance the alignment capabilities of multimodal large language models. The fly in the ointment is that AlignGPT-13B does not perform as well as Qwen-VL on the TextVQA dataset. This may stem from the fact that TextVQA is a text-centric QA task, as it requires identifying text in images to answer questions. AlignGPT is tailored to boost multimodal alignment and might not exhibit strong results in text-focused scenarios.
MLLM-oriented Multi-modal Benchmarks.
We apply AlignGPT to seven recent popular multimodal benchmarks, as shown in Table 3.3. We discover that, apart from LLaVA-1.5-13B, AlignGPT-7B surpassed all previous multimodal models. This shows that our model has strong generalization ability. Additionally, compared to LLaVA-1.5-13B, AlignGPT-13B has shown improvements on most datasets, particularly achieving good advancements on the MME, MMB, and LLaVAW benchamrks. This further validates the efficacy of both global and local alignment capabilities.
Table 3: Influence of different number of alignment level.
{tabu}
l|c|ccccccccc Method Alignment Level VQAv2 GQA VisWiz SQAI TextVQA POPE MME MMB SEEDI
AlignGPT Number=4 79.0 62.9 52.3 68.7 58.3 86.2 1463.8 67.2 66.5
AlignGPT Number=6 79.0 62.7 51.2 68.9 58.3 85.8 1436.3 67.3 66.2
AlignGPT Number=8 79.1 62.9 54.2 68.5 58.4 86.0 1527.4 67.3 66.5
AlignGPT Number=10 79.1 62.6 53.0 67.8 58.4 86.2 1481.4 66.4 66.7
Table 4: Influence of local and global alignment.
{tabu}
c|ccc|cccccccccc Settings Average Local Global VQAv2 GQA VisWiz SQAI TextVQA POPE MME MMB SEEDI
(a) ✘ ✔ ✘ 79.1 62.7 53.3 67.9 58.6 85.9 1467.1 66.9 66.3
(b) ✘ ✘ ✔ 79.1 62.9 52.6 68.3 58.4 85.9 1502.9 66.3 66.2
(c) ✔ ✘ ✔ 79.0 62.8 52.5 68.6 58.4 85.6 1492.5 67.0 66.0
(d) ✘ ✔ ✔ 79.1 62.9 54.2 68.5 58.4 86.0 1527.4 67.3 66.5
5.2 Ablation Study
Without loss of generality, we choose AlignGPT-7b for the ablation study to investigate the effects of different components.
Impact of Number of Alignment Levels.
To investigate the effect of the number of alignment levels $N$ on AlignGPT, we vary the value of $N$ in the range of [4, 10] with a step size of 2. Table  5.1 shows the performance of AlignGPT with different $N$ on nine datasets. Actually, AlignGPT can achieve good results at $N=4$, and their performance remains stable as the number of alignment levels increases. Depending on the trajectory of the curve, their performance has an initial upward trend and then flattens out. These observations indicate that AlignGPT can improve the alignment capabilities of multi-modal large language models based on a small number of alignment levels. Finally, according to the trend of the curve, we set $N$ to 8.
Impact of Local and Global Alignment.
During the instruction-tuning phase, we assign global alignment capabilities and local alignment capabilities to the instructions of each task. Here, we explore the role of separate global alignment capabilities and local alignment capabilities on AlignGPT. Among them, “Local” refers to the local alignment capabilities derived by assigning different weights to various local alignment vectors using a gate network. “Global” denotes the global alignment capabilities, and “Average” represents the local alignment capabilities obtained by assigning equal weights to each local alignment vector. The performance of these four strategies (settings a-d) is presented in Table 5.1. As we can see, setting(a) and setting(b) demonstrate divergent performances in downstream tasks, which can be attributed to the different demands these tasks place on global and local alignment capabilities. It is worth noting that setting (a) and setting (b) perform worse than our final approach (setting d) on most datasets, which verifies the necessity and superiority of the combination of global and local alignment capabilities. Additionally, the performance of setting (c) is inferior to that of setting (d), a possible reason being that the demands for local alignment capabilities in each downstream task are dynamically variable.
Table 5: Influence of different input image resolutions.
{tabu}
l|l|c|ccccccc Method LLM Resolution VQAv2 GQA SQAI TextVQA POPE MMB SEEDI
AlignGPT Vicuna-7B 336 79.1 62.9 68.5 58.4 86.0 67.3 66.5
AlignGPT Vicuna-7B 672 79.7 63.3 68.3 60.3 86.8 67.2 66.5
AlignGPT Vicuna-7B 1008 79.8 63.4 68.2 60.3 86.8 67.2 66.6
Table 6: Influence of different large language models.
{tabu}
l|l|c|ccccccc Method LLM Resolution VQAv2 GQA SQAI MME MMB MMBCN SEEDI
AlignGPT LLaMA2-7B-Chat 336 79.1 62.9 65.9 1500.8 66.6 57.9 66.4
AlignGPT Vicuna-7B 336 79.1 62.9 68.5 1527.4 67.3 59.9 66.5
AlignGPT LLaMA3-8B-Base 336 79.6 63.1 70.4 1539.7 72.0 67.7 68.2
5.3 Discussion
Impact of different input image resolutions.
Image resolution plays a crucial role in vision-language tasks as higher resolutions help reduce image blurring and enhance the understanding of image-text alignment. To evaluate the impact of resolution changes on the performance of multimodal tasks, we increase the image resolution from 336 to 1008, with the resulting performance changes detailed in Table 5.2. The study results show that higher image resolutions can improve model performance on most multimodal tasks. For example, the score for VQAv2 increased from 79.1 to 79.8, while the score for TextVQA rose from 58.4 to 60.3. Meanwhile, the performance of the POPE improve by 0.8. These results highlight that appropriately increasing image resolution is an effective strategy for enhancing performance in studies of multimodal large language models.
Impact of different large language models.
We also explore the impact of the large language model on the performance of AlignGPT, specifically testing three models: LLaMA-2-7B-Chat, Vicuna-v1.5-7B, and the latest LLaMA-3-8B-Base. The results are shown in Table 5.2. Initially, we observe that LLaMA-3-8B-Base achieves the best performance, followed by Vicuna-v1.5-7B, with LLaMA-2-7B-Chat performing the worst, which is reasonable given LLaMA-3-8B-Base’s larger parameter size and richer training data. Besides, we find that compared to the VQA tasks, Vicuna-v1.5-7B performs better on multimodal benchmarks such as MME, MMB, and SEEDI than LLaMA-2-7B-Chat, possibly because Vicuna-v1.5-7B underwent supervised instruction-tuning with ShareGPT data, and ShareGPT contains some background knowledge related to downstream tasks.
5.4 Qualitative Results
Figure 3 presents a comparative analysis of our model with MiniGPT-v2 DBLP:journals/corr/abs-2310-09478 and LLaVA-1.5 DBLP:journals/corr/abs-2310-03744 . When a user submits an image alongside the instruction “Can you see the bright blue feathers on the bird’s wing?”, MiniGPT-v2 and LLaVA-1.5 both return an incorrect answer “Yes”. In contrast, our model produces accurate result “No”, thereby demonstrating that AlignGPT can effectively enhance the model’s alignment capability. In Figure 4, we further demonstrate the responses of AlignGPT under different levels of alignment capability. We find that with lower alignment levels, the model may only focus on certain regions of the image, resulting in an undercount of the total number of apples; whereas with higher alignment levels, the model considers the entire image area, thus achieving accurate apple quantity estimation. This finding once again underscores the necessity of enhancing the alignment capability of MLLMs.
(a) MiniGPT-v2
(b) LLaVA-1.5
(c) AlignGPT
Figure 3: Comparison with other models.
(a) alignment level=1
(b) alignment level=4
(c) alignment level=7
Figure 4: The responses of AlignGPT under different levels of alignment capability.
6 Conclusion
In this paper, we propose AlignGPT, a novel multimodal large language model designed to bolster the alignment capabilities of MLLMs. Our approach involves utilizing the alignment level of data as a control signal during pre-training to effectively handle the varying degrees of alignment in image-text pairs. Subsequently, in the instruction-tuning phase, we begin by exploiting these control signals to shape different levels of alignment capabilities. Continuing from this, we go beyond assigning global alignment capabilities to instructions of each task; we also dynamically configure distinct local alignment capabilities based on the specific demands of each instruction. Results from numerous experiments indicate that our AlignGPT achieves better performance than other state-of-the-art MLLMs. Further analysis validates the superiority of our AlignGPT.